# HERBench: A Benchmark for Multi-Evidence Integration in Video Question Answering

### v1

@@ Abstract
Video Large Language Models (Video-LLMs) are rapidly improving, yet current Video Question Answering (VideoQA) benchmarks often allow questions to be answered from a single salient cue, under-testing reasoning that must aggregate multiple, temporally separated visual evidence. In this direction, we present HERBench, a VideoQA benchmark purpose-built to assess multi-evidence integration across time. Each question is constructed to require aggregating at least three non-overlapping evidential cues across distinct video segments (so neither language priors nor a single snapshot can suffice). HERBench comprises 26K five-way multiple-choice questions organized into twelve compositional tasks that probe identity binding, cross-entity relations, temporal ordering, co-occurrence verification, and counting. To make evidential demand measurable, we introduce the Minimum Required Frame-Set (MRFS)-the smallest number of frames a model must fuse to answer correctly-and show that HERBench imposes substantially higher demand than prior datasets (mean MRFS $5.5$ vs. 2.62.6-$4.2$). Evaluating 13 state-of-the-art Video-LLMs on HERBench reveals pervasive failures: accuracies of 31–42% are only slightly above the 20% random-guess baseline. We disentangle this failure into two critical bottlenecks: (1) a retrieval deficit, where frame selectors overlook key evidence, and (2) a fusion deficit, where models fail to integrate information even when all necessary evidence is provided. By making cross-time evidence both unavoidable and quantifiable, HERBench establishes a principled target for advancing robust, compositional video understanding. [Uncaptioned image]
Figure 1: From Single-Cue to Multi-Evidence Integration. While existing benchmarks like MVBench [li2024mvbench] (top) often focus on short-term attributes solvable via single salient frames or language priors, HERBench (bottom) enforces a high Evidential Requirement (ER). In this Temporal Shot Ordering example, the model must identify and temporally bind four distinct, non-overlapping visual evidence dispersed across the video to reconstruct the correct sequence. This design ensures that successful answering requires genuine multi-evidence integration rather than reliance on static shortcuts.
@@ 1 Introduction
As Video Large Language Models [qwen3vl_8b_2025, llava_onevision15_8b_2025, ovis25_9b_2025] achieve strong scores on established VideoQA benchmarks [lei2018tvqa, jang2017tgifqa, xiao2021nextqa, tapaswi2016movieqa, mangalam2023egoschema, fu2024videomme, longvideobench2024], their video understanding capabilities appear to be rapidly emerging. However, recent audits reveal these high scores often stem from language priors or single-cue shortcuts rather than grounded temporal reasoning [tempcompass2024, xiao2024visground], causing models to fail tasks that explicitly require multi-hop inference [girdhar2020cater, yi2020clevrer, grundemclaughlin2021agqa]. In contrast, tasks like Referring Video Object Segmentation (RVOS) demonstrate that robust, multi-frame aggregation is achievable, as models successfully link instances across occlusions and appearance changes [Gavrilyuk_2018_CVPR, Seo_2020_ECCV, Botach_2022_CVPR, Wu_2022_CVPR].
We advocate centering evaluation on evidential requirement, because single-cue questions fail to measure multi-evidence integration. We define the Evidential Requirement (ER) as the minimum number of distinct, non-redundant visual evidence needed for an answer. High-ER items make compositional reasoning, such as temporal binding and clue combination, unavoidable [xiao2021nextqa, xiao2021star, yi2020clevrer]. Controlling ER therefore distinguishes models that integrate information from those that rely on isolated cues [tempcompass2024]. This approach makes aggregation measurable, aligns VideoQA with real-world reasoning, and offers a principled path for progress beyond single-cue success [xiao2021nextqa, xiao2021star, yi2020clevrer].
We introduce HERBench (High Evidential Requirement Benchmark), where questions across twelve compositional subtasks (e.g., entity binding, temporal ordering) are constructed to structurally enforce $k\geq 3$ distinct pieces of evidence, as presented in Figure 1. To measure this, we present the Minimum Required Frame-Set (MRFS) metric, defined as the minimum number of frames needed for a correct answer. Cross-benchmark validation confirms our high-ER design: HERBench’s mean MRFS of 5.5 far exceeds the 2.6 to 4.2 observed in existing benchmarks and enables principled, ER-focused diagnostics.
Our evaluation of state-of-the-art Video-LLMs exposes two critical bottlenecks. Finding 1: Frame selection is a major bottleneck. While adaptive selectors [tang2025aks, liu2025bolt] outperform uniform sampling, they still lag significantly behind ground-truth keyframes. Finding 2: Multi-evidence reasoning is also a bottleneck. Even with ground-truth frames, models achieve only modest accuracy because they fail to assign proper importance to all critical frames and struggle to integrate them. Progress therefore requires advances in both frame selection and multi-evidence reasoning.
@@
  • We propose the Minimum Required Frame-Set (MRFS) metric: a measure of the smallest number of frames a model must aggregate to answer a question correctly, thereby enabling apples-to-apples comparison across benchmarks (existing benchmarks range from 2.6 to 4.2, whereas HERBench averages 5.5) and powering ER-focused diagnostics.
@@ 2 Related Work
Video-LLMs architectures have evolved from simple feature pooling [damonlpsg2023videollama, Maaz2023VideoChatGPT] to sophisticated systems employing advanced alignment modules (e.g., Q-Formers) and large-scale instruction tuning [li2023blip2, dai2023instructblip, llava_onevision_7b_2024] to bridge the modality gap. Despite massive token capacities in proprietary models like Gemini 2.5 [gemini25_2025] and GPT-4o [openai2024gpt4o], recent audits [tempcompass2024, breaking-down] reveal a persistent failure in robust temporal aggregation. Instead of performing multi-hop inference, these models frequently default to language priors or single-frame shortcuts to solve tasks.
@@
VideoQA evaluation has progressed from short-clip recognition [xu2017video, xu2016msrvtt] to long-form reasoning. While MVBench [li2024mvbench] introduced diverse temporal tasks, its short duration limits long-horizon assessment. Successors like EgoSchema [mangalam2023egoschema], LongVideoBench [longvideobench2024], and Video-MME [fu2024videomme] address this by targeting minute-to-hour scale videos and diverse modalities. However, “long context” does not equate to “high evidential requirement.” Critiques [breaking-down, mangalam2023egoschema] indicate these benchmarks often remain solvable via single salient keyframes, failing to measure genuine multi-step reasoning. HERBench distinguishes itself by explicitly controlling the Evidential Requirement (ER), constructing questions that structurally enforce aggregating multiple ($k\geq 3$) distinct, temporally separated visual cues to assess compositional reasoning rather than memory capacity.
@@ 3 HERBench: High Evidential Requirement Benchmark
In this section, we present HERBench, a VideoQA benchmark explicitly designed to evaluate multi-evidence integration. The section is organized as follows: Sec. 3.1 introduces the twelve tasks grouped into four reasoning families, together with the specific capabilities they target; Sec. 3.2 describes the construction pipeline that enforces a high Evidential Requirement (ER) while suppressing shortcuts; Sec. 3.3 summarizes the dataset scale and corpus statistics; and Sec. 3.4 formalizes ER through the Minimum Required Frame-Set (MRFS) metric and presents the standardized evaluation protocol.
@@ 3.1 Task Taxonomy
To evaluate whether models truly use evidence rather than guess from the most prominent cue, we organize twelve tasks into four families (Figure 2). Each task is constructed so that a correct answer requires aggregating at least $k\geq 3$ distinct observations (ER), enforced via identity binding, set-level aggregation, temporal ordering, or video-wide verification.
@@
These tasks require understanding event order, co-occurrence, and durations, compiling distributed cues into a linear chronology. The three tasks are: 1) [TSO] Temporal Shot Ordering: Arrange four shot descriptions from a trailer into the correct chronological order, using only content cues to reconstruct high-level scene transitions. 2) [MPDR] Multi-Person Duration Reasoning: Compare interval statistics for appearance-described people (e.g., who stayed in view the longest, or who entered/exited first), focusing on fine-grained time-span contrasts across individuals. 3) [ASII] Action Sequence Integrity & Identification: Select the correct ordering of five narrated actions among plausible permutations, stressing micro-level task sequencing rather than scene-level ordering. The ER is driven by ordering and interval comparisons across at least three temporally separated observations, but each task probes a distinct temporal structure.
@@
This family tests binding a uniquely appearance-described target across time to reason about trajectory-dependent properties. Models must maintain a stable reference as the target interacts with the scene. The tasks are: 1) [AGBI] Appearance-Grounded Behavior Interactions: Identify who accompanies or interacts with the target during traversal, emphasizing social and relational cues. 2) [AGAR] Appearance-Grounded Attribute Recognition: Track the target to read out attributes anchored to their immediate local context (e.g., a passerby’s jacket color), focusing on moment-specific attribute extraction. 3) [AGLT] Appearance-Grounded Localization Trajectory: Recover path endpoints and coarse trajectory (e.g., exit method), highlighting global, path-level motion reasoning. This enforces $k\geq 3$ through identity maintenance across separated glimpses, each task centering on a different aspect of target evolution.
@@ 3.2 Benchmark Construction
Figure 3: HERBench Data Construction Pipeline. We employ a tripartite pipeline. (Left) Videos are processed through three parallel streams: 1) Object Tracking and Trajectory Analysis (via RF-DETR and DeepSORT) to produce targets to generate disentangled Appearance (A) and Behavior (B) cards; 2) Shot Segmentation using shot detection with an MLLM description for producing scene descriptions; and 3) Ground Truth Integration refining human verified raw event logs. (Middle) These refined data input are controlled via a Manual Review and then input into an Oriented Task Programming module that programmatically compiles the 12 compositional tasks. (Right) The pipeline enforces rigorous quality control through expert Manual Review and a Text-Only Filtering stage to eliminate language priors, ensuring all final Multiple Choice Questions (MCQs) enforce multi-evidence integration.
@@
This first stream anchors tasks in continuous object dynamics. We employ RF-DETR [zhu2021deformable] and DeepSORT [wojke2017simple] to obtain entity tracks, retaining top entities via a TrackRank score, a composite score favoring appearance rarity (HSV/LBP), trajectory length and frame coverage (see Supplementary). For each track, we generate strictly non-overlapping A-cards (appearance) and B-cards (behavior/trajectory). This decorrelation intentionally separates the identifying appearance from the queried behavior, often placing them in temporally distant frames. This scaffold supports tasks requiring fine-grained interaction and motion analysis:
@@
Finally, this stream moves beyond automated analysis to leverage human verified narrated events [perrett2025hdepic]:
@@
We employ a multi-stage verification protocol governing both the refined data input and the final output. First, the structured cards undergo component-level verification. To ensure referring tasks rely on tracking rather than description matching, we enforce strict disentanglement between A-cards and B-cards via token-level Jaccard similarity checks and manual leakage review. The validated components are processed by the Oriented Task Programming module to instantiate the tasks. Next, to explicitly suppress language priors, we apply a Text-Only Filtering stage: we discard any question correctly answered by $\geq 3$ of 4 blind LLMs (Qwen2-7B [qwen2], Qwen2.5-7B [qwen2_5], Llama-3-8B [llama3], and Vicuna-7B v1.5 [vicuna15]). Finally, experts conduct a verification on a stratified 15% sample to audit constraint satisfaction. This stage specifically targets the minimum multi-frame requirement ($k\geq 3$), rejecting items solvable by single frames ($\sim 18\%$ rejection rate) or those lacking unique, objective ground-truth answers.
@@ 3.3 Benchmark Statistics
Scale and Scope.
HERBench contains 26,806 multiple-choice questions sampled from 336 unique videos. These items are organized into 12 compositional tasks across four categories (Sec. 3.1). Frequent query terms are visualized in Figure 4 (left).
Video Corpus Characteristics.
The corpus is curated to test reasoning over substantial durations (avg. 395s, range 60–2100s), ensuring that required evidential cues are temporally dispersed. It covers balanced environments (46% indoor, 28% outdoor, 26% mixed) and diverse perspectives, including egocentric, surveillance, and cinematic views. Videos are sourced from HD-EPIC [perrett2025hdepic], WildTrack [wildtrack], PersonPath22 [personpath22], and publicly available movie trailers on YouTube; source distribution appears in Figure 4 (center).
Question Format and Density.
Questions utilize a five-way multiple-choice format, establishing a uniform 20% random-guess baseline. Crucially, correct answers are statistically balanced across options (A-E) to eliminate positional bias. Figure 4 (right) details the question distribution per task.
@@ 3.4 Evidential Requirement & the MRFS Metric
Answering a VideoQA item may require fusing evidence from multiple, temporally separated moments, or it may be solvable from a single salient frame. To make this requirement measurable, we introduce the Minimum Required Frame-Set (MRFS): the smallest number of frames that must be fused for a model to answer a given question correctly. Aggregating MRFS over items yields a benchmark-level statistic that quantifies evidential demand, higher mean MRFS indicates that questions cannot be solved by single-cue shortcuts and instead require multi-moment integration.
@@
Benchmark Questions Video length (s) MRFS
NExT-QA 4,996 44 2.61
MVBench 4,000 16 3.52
LongVideoBench 6,678 473 4.07
HERBench 26,806 395 5.49
@@
MRFS measures across benchmarks.
Because MRFS is defined with respect to $(f,r,x)$, cross-benchmark comparability requires fixing these components. We standardize on $f=$ Qwen2.5-VL [qwen25vl_72b], $r=$ AKS adaptive keyframe sampling [tang2025aks], and $x=16$ frames. This protocol isolates the dataset’s evidential requirement from model or selector variability. As seen in Table 1, HERBench achieves a mean MRFS of 5.49, a $\sim$35% increase over the next-highest benchmark, LongVideoBench (4.07), and far exceeds MVBench (3.52) and NExT-QA (2.61). Notably, this high ER is achieved with a shorter average video duration (395s) than LongVideoBench (473s), indicating the difficulty arises from evidential density, not just length. These results confirm that HERBench makes multi-evidence integration unavoidable. MRFS turns an otherwise qualitative notion like “how much evidence does a question require?” into a quantitative, model- and selector-standardized measurement. Reporting mean MRFS alongside accuracy thus separates success via single-cue shortcuts from genuine multi-evidence reasoning and makes benchmark evidential demand explicit.

### v2

@@ Abstract
Video Large Language Models (Video-LLMs) are improving rapidly, yet current Video Question Answering (VideoQA) benchmarks often admit single-cue shortcuts, under-testing reasoning that must integrate evidence across time. We introduce HERBench, a benchmark designed to make multi-evidence integration unavoidable: each question requires at least three non-overlapping cues drawn from distinct video segments. HERBench contains 26,806 five-way multiple-choice questions across 12 compositional tasks. To make evidential demand measurable, we introduce the Minimum Required Frame-Set (MRFS), the smallest number of frames a model must fuse to answer correctly, and show that HERBench imposes higher evidential demand than prior benchmarks. Evaluating 13 state-of-the-art Video-LLMs yields only 31–42% accuracy, only modestly above the 20% random-guess baseline. We disentangle this failure into two critical bottlenecks: (1) a retrieval deficit, where frame selectors overlook key evidence, and (2) a fusion deficit, where models fail to integrate information even when all necessary evidence is provided. HERBench thus provides a principled benchmark for studying robust multi-evidence video understanding. [Uncaptioned image]
Figure 1: From Single-Cue to Multi-Evidence Integration. While existing benchmarks like MVBench [22] (top) often focus on short-term attributes solvable via single salient frames or language priors, HERBench (bottom) enforces a high Evidential Requirement (ER). In this Temporal Shot Ordering example, the model must identify and temporally bind four distinct, non-overlapping visual evidence dispersed across the video to reconstruct the correct sequence. This design ensures that successful answering requires genuine multi-evidence integration rather than reliance on static shortcuts.
@@ 1 Introduction
As Video Large Language Models [2, 1, 27] achieve strong scores on established VideoQA benchmarks [18, 16, 50, 44, 29, 11, 48], their video understanding capabilities appear to be rapidly emerging. However, recent audits reveal these high scores often stem from language priors or single-cue shortcuts rather than grounded temporal reasoning [26, 51], causing models to fail tasks that explicitly require multi-hop inference [14, 55, 15]. In contrast, tasks like Referring Video Object Segmentation (RVOS) demonstrate that robust, multi-frame aggregation is achievable, as models successfully link instances across occlusions and appearance changes [12, 39, 4, 49].
We advocate centering evaluation on evidential requirement, because single-cue questions fail to measure multi-evidence integration. We define the Evidential Requirement (ER) as the minimum number of distinct, non-redundant visual evidence needed for an answer. High-ER items make compositional reasoning, such as temporal binding and clue combination, unavoidable [50, 47, 55]. Controlling ER therefore distinguishes models that integrate information from those that rely on isolated cues [26]. This approach makes aggregation measurable, aligns VideoQA with real-world reasoning, and offers a principled path for progress beyond single-cue success [50, 47, 55].
We introduce HERBench (High Evidential Requirement Benchmark), where questions across twelve compositional subtasks (e.g., entity binding, temporal ordering) are constructed to structurally enforce $k\geq 3$ distinct pieces of evidence, as presented in Figure 1. To measure this, we present the Minimum Required Frame-Set (MRFS) metric, defined as the minimum number of frames needed for a correct answer. Cross-benchmark comparison under a canonical MRFS protocol confirms our high-ER design: HERBench attains the highest mean MRFS among the benchmarks considered and enables principled ER-focused diagnostics.
Our evaluation of state-of-the-art Video-LLMs exposes two critical bottlenecks. Finding 1: Frame selection is a major bottleneck. While adaptive selectors [43, 25] outperform uniform sampling, they still lag significantly behind ground-truth keyframes. Finding 2: Multi-evidence reasoning is also a bottleneck. Even with ground-truth frames, models achieve only modest accuracy because they fail to assign proper importance to all critical frames and struggle to integrate them. Progress therefore requires advances in both frame selection and multi-evidence reasoning.
@@
  • We propose the Minimum Required Frame-Set (MRFS) metric: a measure of the smallest number of frames a model must aggregate to answer a question correctly, thereby enabling apples-to-apples comparison across benchmarks and powering ER-focused diagnostics.
@@ 2 Related Work
Video-LLMs architectures have evolved from simple feature pooling [57, 28] to sophisticated systems employing advanced alignment modules (e.g., Q-Formers) and large-scale instruction tuning [20, 8, 19] to bridge the modality gap. Despite massive token capacities in proprietary models like Gemini 2.5 [7] and GPT-4o [33], recent audits [26, 10] reveal a persistent failure in robust temporal aggregation. Instead of performing multi-hop inference, these models frequently default to language priors or single-frame shortcuts to solve tasks.
@@
VideoQA benchmarks have evolved from short-clip recognition [52, 53] toward longer-form evaluation, but they often probe different aspects of video understanding. MVBench [22] broadened the task space with diverse temporal questions, yet its short clips limit the assessment of long-horizon assessment. More recent benchmarks, including EgoSchema [29], LongVideoBench [48], and Video-MME [11], expand temporal scope, while MINERVA [32] emphasizes complex multi-step reasoning and reasoning-trace auditing. HERBench targets a different axis: not only how long the context is, but how many visual pieces of evidence must be combined. While phenomena such as temporal ordering and counting appear in prior benchmarks, they are typically not formulated so that solving them requires aggregating multiple temporally separated cues. HERBench instead explicitly controls the Evidential Requirement (ER): each question is constructed to require at least three non-overlapping cues drawn from distinct moments in the video, and the benchmark’s oracle-frame design enables retrieval failures to be disentangled from fusion failures.
@@ 3.1 Task Taxonomy
To evaluate whether models truly integrate evidence rather than rely on a salient cue, we organize 12 tasks into four reasoning families (Figure 2). These families recast familiar VideoQA settings under structural $k\geq 3$ constraints over long videos, while also introducing tasks for appearance-grounded identity binding, set-level identity maintenance, and region-conditioned aggregation.
@@
These tasks require understanding event order, co-occurrence, and durations, compiling distributed cues into a linear chronology. The three tasks are: 1) [TSO] Temporal Shot Ordering: Arrange four shot descriptions from a trailer into the correct chronological order, using only content cues. 2) [MPDR] Multi-Person Duration Reasoning: Compare interval statistics for appearance-described people (e.g., who stayed in view the longest, or who entered/exited first), focusing on fine-grained time-span contrasts across individuals. 3) [ASII] Action Sequence Integrity & Identification: Select the correct ordering of five narrated actions among plausible permutations, stressing micro-level task sequencing rather than scene-level ordering. The ER is driven by ordering and interval comparisons across at least three temporally separated observations, but each task probes a distinct temporal structure.
@@
This family tests binding a uniquely appearance-described target across time to reason about trajectory-dependent properties. Models must maintain a stable reference as the target interacts with the scene. The tasks are: 1) [AGBI] Appearance-Grounded Behavior Interactions: Identify who accompanies or interacts with the target during traversal, emphasizing social and relational cues. 2) [AGAR] Appearance-Grounded Attribute Recognition: Track the target to read out attributes anchored to their immediate local context (e.g., a passerby’s jacket color), focusing on moment-specific attribute extraction. 3) [AGLT] Appearance-Grounded Localization Trajectory: Recover path endpoints and coarse trajectory (e.g., exit method), highlighting global, path-level motion reasoning. This enforces $k\geq 3$ through identity maintenance across separated glimpses, as the target description is composed of cues scattered across distinct moments that must be jointly resolved, with each task centering on a different aspect of target evolution.
@@ 3.2 Benchmark Construction
Figure 3: HERBench construction pipeline. Videos are processed through three streams—object tracking and trajectory analysis, shot segmentation, and ground-truth integration (refining human verified raw event logs) —whose outputs are compiled by oriented task programming and filtered by manual review and text-only bias checks.
@@
This first stream anchors tasks in continuous object dynamics. We employ RF-DETR [38] and DeepSORT [46] to obtain entity tracks, retaining top 20% entities via a TrackRank score, a composite score favoring appearance rarity, trajectory length and frame coverage (see Supplementary). For each track, we generate strictly non-overlapping A-cards (appearance) and B-cards (behavior/trajectory). This decorrelation intentionally separates the identifying appearance from the queried behavior, often placing them in temporally distant frames, and enables a new class of appearance-grounded identity-binding tasks that cannot be solved by local attribute lookup alone. Full error quantification and noise control details are provided in the Supplementary.
This scaffold supports tasks requiring fine-grained interaction and motion analysis:
@@
Finally, this stream moves beyond automated analysis to leverage human verified narrated events [35]:
@@
The refined representations are compiled by oriented task programming into multiple-choice questions, then filtered through several safeguards. We enforce A/B-card disentanglement via token-level similarity checks and manual leakage review, discard items solved by $\geq 3$ of 4 blind LLMs, and perform expert verification on a stratified 15% sample to confirm minimum frame-set ($k\geq 3$) compliance and answer uniqueness. Human validation further confirmed both answerability and oracle-evidence quality: annotators achieved 88.8% accuracy with full-video access and 95.7% in an oracle-frame setting; details are deferred to the Supplementary Material. As HERBench is partially instantiated through an oriented task-programming pipeline, some residual systematic artifacts may nevertheless remain.
@@ 3.3 Benchmark Statistics
HERBench contains 26,806 five-way multiple-choice questions from 336 unique videos spanning 12 tasks and four reasoning families. The videos are long-form (395 s on average; range 60–2100 s), diverse in source and viewpoint, and chosen to ensure temporal dispersion of evidence; Figure 4 summarizes the resulting vocabulary, source distribution, and per-task question counts.
@@ 3.4 Evidential Requirement & the MRFS Metric
VideoQA questions may require fusing distributed evidence or be solvable via single salient frames. To quantify this, we introduce the Minimum Required Frame-Set (MRFS): the smallest number of frames a model must fuse to answer correctly. A higher mean MRFS confirms that questions resist single-cue shortcuts and genuinely demand multi-moment integration. Although prior metrics capture important temporal dynamics, they measure fundamentally different properties. Temporal Indispensability (1-frame vs. N-frame performance) [9] tests for static shortcuts, while Certificate Length [11, 29] measures the human-annotated temporal span needed for verification. In contrast, MRFS is an automated, model-centric metric that isolates the exact number of frames a model must fuse, directly quantifying the multi-evidence aggregation challenge without relying on human annotation.
@@
Benchmark
# Videos
# Questions
@@
Lang.
debias
Enforced
fusion
Absence
check
Oracle
frames
TemporalBench 2,179 9,867 2.21
MMBench-Video 609 1,998 4.41
Video-MME 900 2,700 5.31
AGQA (balanced) 9,600 3.9 M 3.42
CVRR-ES 217 2,400 2.77
LongVideoBench 3,763 6,678 4.07
NExT-QA 5,440 99,736 2.61
MINERVA 223 1,515 5.14
MVBench 4,000 4,000 3.52
HERBench 336 26,806 5.49
Table 1: Benchmark comparison under the canonical MRFS protocol. MRFS is reported with $f=$ Qwen2.5-VL, $r=$ AKS, and $x=16$. Enforced fusion indicates whether correct answering structurally requires combining multiple temporally separated visual cues.
Cross-benchmark MRFS comparison.
To enable cross-benchmark comparison, we fix a canonical MRFS protocol: Qwen2.5-VL as backbone, AKS as frame selector, and $x=16$ frames. Table 1 compares HERBench to prior benchmarks under this shared setting, together with complementary design properties. HERBench achieves the highest mean MRFS (5.49) in the comparison and is uniquely defined by the combination of language debiasing, structurally required temporal fusion, absence checking, and oracle-frame availability. Among prior benchmarks, MINERVA [32] is the closest in spirit, pairing language debiasing with a relatively high MRFS (5.14), but focusing on multi-step reasoning and reasoning-trace auditing rather than enforced multi-frame evidence aggregation. The benchmark ordering remains stable across alternative backbones and selectors (see Supplementary), and the comparison with LongVideoBench [48] suggests that HERBench derives its difficulty from evidential density rather than duration alone.